OpenStudy (anonymous):

Given the following functions f(x)=x^2-3x+4 and g(x)=x-2 fine (f0g) (x)

OpenStudy (anonymous):

kizzy please take a look at your previous post

OpenStudy (anonymous):

\[f(g(x))=f(x-2)=(x-2)^2-3(x-2)+4\]

OpenStudy (anonymous):

hey satellite i need your help

OpenStudy (anonymous):

then some algebra \[(x-2)^2-3(x-2)+4=x^2-4x+4-3x+6+4\] \[=x^2-7x+14\]

OpenStudy (anonymous):

what do you need hikari?

OpenStudy (anonymous):

go to a is about and there a link